The sum of two numbers is 80. If one is 20 greater than the other, find the numbers
Found 2 solutions by mananth, Cromlix:
Answer by mananth(16946)

let the numbers be x & y x>y
x+y=80
x-y=20
add the equations
2x=100
x=50
plug x=50 in the first equation
x+y=80
50+y=80
y=80-50
y=30
The numbers are 50 & 30

Answer by Cromlix(4381)

Small number = x
Larger number = x + 20
x + (x + 20) = 80
x + x + 20 = 80
2x = 80 - 20
2x = 60
x = 30.
Small number = 30
Larger number = 50
Hope this helps.
:-)